# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented Large Language Models

### v4

@@
Abstract.
Recently, code-oriented large language models (Code LLMs) have been widely and successfully exploited to simplify and facilitate programming. With these tools, developers can easily generate the desired complete functional code based on incomplete code snippets and natural language prompts. Unfortunately, a few pioneering works revealed that these Code LLMs are vulnerable to backdoor and adversarial attacks. The former poisons the training data or model parameters, hijacking the LLMs to generate malicious code snippets when encountering the trigger. The latter crafts malicious adversarial input codes to reduce the quality of the generated codes. However, both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes.
@@ Abstract.
This paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA). TPIA generates non-functional perturbations containing the information of malicious instructions and inserts them into the victim’s code context by spreading them into potentially used dependencies (e.g., packages or RAG’s knowledge base). It induces the Code LLMs to generate attacker-specified malicious code snippets at the target location. In general, we compress the attacker-specified malicious objective into the perturbation by adversarial optimization based on greedy token search. We collect 13 representative malicious objectives to design 31 threat cases for three popular programming languages. We show that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation. Our work alerts a new practical threat of using Code LLMs.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Recently, generative artificial intelligence (GenAI) techniques have been widely and successfully exploited in code engineering, such as automated programming and code generation. In particular, code-oriented large language models (Code LLMs), e.g., CodeX (Wojciech Zaremba, 2021), CodeGemma (Team et al., 2024), and CodeGeex (Zheng et al., 2023), have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus (Zhang et al., 2023; Li et al., 2023b; Ding et al., 2024) and the fundamental paradigm in commercial practice (Friedman, 2021; cod, 2024).. They can efficiently generate complete functional code based on incomplete code snippets and text prompts. Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot (Friedman, 2021) developed jointly by Github (Git, 2024) and OpenAI (Ope, 2024). These applications can provide real-time code completion suggestions in the Integrated Development Environment (IDE), significantly improving programming efficiency.
@@
Despite their remarkable performance, recent studies revealed that Code LLMs are vulnerable to some well-designed attacks, which maliciously manipulate the functional execution of Code LLMs in practice. In general, existing attacks can be divided into two main categories. (1) Backdoor attacks (Schuster et al., 2021; Li et al., 2023a; Yang et al., 2024). They intend to compromise the models by poisoning the training data (Schuster et al., 2021; Yang et al., 2024) or model parameters (Li et al., 2023a), such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns. They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications (Li et al., 2023a), or implementation of less secure encryption rules or protocols (Schuster et al., 2021). (2) Adversarial attacks (Zhou et al., 2022; Bielik and Vechev, 2020; Jha and Reddy, 2023). They aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.(Jha and Reddy, 2023) slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score (Ren et al., 2020).
@@
However, we reveal that existing attacks suffer from some inherent limitations, hindering their realistic threats to some extent. Specifically, backdoor attacks can achieve specific malicious objectives but require the capability of interfering the model training, e.g., modifying some training samples or even model parameters. In real-world scenarios, developers of Code LLMs or integrated applications usually obtain the models or datasets from trusted entities (e.g., renowned research institutes or companies). This mitigates the backdoor threats from the source. Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult. Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks (Bielik and Vechev, 2020; Zhou et al., 2022; Jha and Reddy, 2023) can hardly achieve precise malicious objectives, although they do not require intervention in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, can inevitably weaken Code LLMs’ understanding of the input. Driven by these findings, an intriguing and critical question arises: Has the application of Code LLMs already achieved a sufficient level of security in practice?
@@
Unfortunately, the answer to the above question is negative. In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger (Wallace et al., 2019) and Greedy Coordinate Gradient (GCG) (Zou et al., 2023). Please find more details in Section 2 and Appendix A.1.. TPIA leads to a more realistic threat by inheriting the advantageous features of both backdoor and adversarial attacks. In general, as shown in Figure 1, the TPIA attacker generates a concise non-functional perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion) , and inserts it into the victim’s code context. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations. Arguably, attackers can easily introduce particular malicious non-functional perturbations to the victim’s code context in practice. For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs (Chen et al., 2024) via retrieval-augmented generation (Lewis et al., 2020). In particular, different from supply chain attacks (Ohm et al., 2020) that can be easily detected by static analysis (Li et al., 2018b) or sandboxing (Alhamdan and Staicu, 2023) since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
@@
Figure 1. The attack paragdim of our TPIA. A concise non-functional attacker-generated perturbation is injected into one of the project’s files. When the victim user completes the code via the Code LLM, the model will be induced by the incorporated attacker-specified code to generate an unsafe output at specified position, e.g., in edit_part() function.
@@
In general, our attack consists of two main stages. In the first stage, we construct the input and target output based on the attacker-specified objective. We represent the malicious objective as an output tuple (target position, target code), which aims to induce the Code LLM to generate the target code at the target position. We further expand the tuple into a complete task code, including a conditional code (e.g., necessary package references), context code (e.g., noises), and position code to simulate the actual attack scenario. In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective. Specifically, we obtain the base Code LLM of the integrated application and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code. This computation involves two important optimizations: (1) Forward reasoning enhancement: it allows the optimization direction of adversarial computation to be closer to the actual reasoning process of the Code LLM (forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. (2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack.
@@
We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) (Top, 2024) across 3 programming languages. There are a total of 31 threat cases since some of these malicious objectives apply only on C/C++. We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b, and CodeGeex2-6b. The results demonstrate that TPIA can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness. In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX (cod, 2024) (with more than 535k installations in the VScode (VSC, 2024) extension shop) and Github Copilot (Friedman, 2021) (developed by Github and OpenAI (Ope, 2024)). These applications correspond to the gray-box and black-box attack scenarios, respectively. Our TPIA achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its practical threat in the real world.
@@
    We conduct extensive experiments to evaluate our TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 widely used commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
2. Background and Related Work
2.1. Code LLMs for Code Completion
@@
Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation (Sun et al., 2022; Roziere et al., 2020), and code summarization (McBurney and McMillan, 2016; Ahmad et al., 2020). Many models have been developed to achieve these tasks, including CodeGen (Nijkamp et al., 2022), CodeBERT (Feng et al., 2020), and CodeT5 (Wang et al., 2021). The emergence of LLMs provides new alternatives to the coding tasks. By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex (Wojciech Zaremba, 2021), CodeLlama (Roziere et al., 2023), CodeGemma (Team et al., 2024) and CodeGeex (Zheng et al., 2023). These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast array of functionally complex code (Xia et al., 2023).
@@
In this paper, we focus on code completion (Li et al., 2023b; Ding et al., 2024; Zhang et al., 2023). This is a paradigmatic task that involves generating new code from incomplete input (with natural language) using Code LLMs, which have been integrated into commercial applications (Friedman, 2021; cod, 2024). However, this development paradigm may encounter some potential threats. On the one hand, Code LLMs can be easily influenced by the code context, which alters the result of code generation. On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs. These threats have already occurred in the real world (Cri, 2025), highlighting the necessity of its security research.
2.2. Backdoor and Adversarial Attacks
@@
Backdoor attacks target the training process of DNNs (Liu et al., 2020; Li et al., 2022; Zhang et al., 2024). These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present. In the code domain, there are generally two strategies to achieve this, including dataset poisoning (Schuster et al., 2021; Yang et al., 2024) and model poisoning (Li et al., 2023a). The former requires the attacker to modify the code dataset used for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM used for application development. However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned research institutes or companies). This mitigates the backdoor threats from the source. Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult. Arguably, backdoors cannot be implanted into a Code LLM without controlling the model training process.
@@
Different from backdoor attacks, adversarial attacks primarily target the model’s inference process (Li et al., 2018a; Doan et al., 2021; Wang et al., 2022). Adversarial attacks induce the vulnerabilities of models by introducing well-crafted perturbations to the input data, leading the model to make incorrect output. In the code domain, existing adversarial attacks (Bielik and Vechev, 2020; Zhou et al., 2022; Jha and Reddy, 2023) aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.(Jha and Reddy, 2023) slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score (Ren et al., 2020). However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input. Accordingly, adversarial attacks are usually difficult to pose a serious threat to Code LLMs in practice.
2.3. Prompt Injection and Indirect Prompt Injection Attacks
@@
Prompt injection (PI) attacks, are a special form of attack targeting LLMs (DAN, 2023; Liu et al., 2023b; Zou et al., 2023). A malicious LLM user aims to bypass the safety guardrails of the LLM and forces it to perform undesired responses, e.g., details of creating bomb. A general way to carry out these attacks is by submitting a carefully designed prompt query to the LLM. In the code domain, existing research (Jiang et al., 2023; Liu et al., 2023a; Fang et al., 2024) focuses on code-oriented attack objectives, such as using LLMs to create malware. In general, these attacks target the developers of models (to make their models unsafe) instead of users, which is out of the scope of this paper.
@@
Indirect prompt injection (IPI) attack (Greshake et al., 2023), on the other hand, considers a different attack scenario than PI. The attacker attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites or providing false information. This attack simultaneously threatens both LLM-integrated applications (by reducing usability) and victim users (by receiving malicious responses). In the code domain, the only existing work (Greshake et al., 2023) makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it needs to insert a large amount of readable comments containing the malicious intent, which can be easily noticed by programmers, as shown in Figure 11. Arguably, the existing IPI attack does not pose a realistic threat to Code LLMs.
3. Problem Statement
3.1. Code Completion Task

### v5

# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs

@@
Abstract
@@
Recently, code-oriented large language models (Code LLMs) have been widely and successfully exploited to simplify and facilitate programming. Unfortunately, a few pioneering works revealed that these Code LLMs are vulnerable to backdoor and adversarial attacks. The former poisons the training data or model parameters, hijacking the LLMs to generate malicious code snippets when encountering the trigger. The latter crafts malicious adversarial input codes to reduce the quality of the generated codes. In this paper, we reveal that both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes. To alleviate these problems, this paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA). TPIA generates non-functional perturbations containing the information of malicious instructions and inserts them into the victim’s code context by spreading them into potentially used dependencies (e.g., packages or RAG’s knowledge base). It induces the Code LLMs to generate attacker-specified malicious code snippets at the target location. In general, we compress the attacker-specified malicious objective into the perturbation by adversarial optimization based on greedy token search. We collect 13 representative malicious objectives to design 31 threat cases for three popular programming languages. We show that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation.
I Introduction
@@
Recently, generative artificial intelligence (GenAI) techniques have been widely and successfully exploited in code engineering, such as automated programming and code generation. In particular, code-oriented large language models (Code LLMs), e.g., CodeX [53], CodeGemma [49], and CodeGeex [61], have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus [60, 34] and the fundamental paradigm in commercial practice [19, 2].. They can efficiently generate complete functional code based on incomplete code snippets and text prompts. Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot [19] developed jointly by Github [3] and OpenAI [4]. These applications can provide real-time code completion suggestions in the Integrated Development Environment (IDE), significantly improving programming efficiency.
@@
Despite their remarkable performance, recent studies revealed that Code LLMs are vulnerable to some well-designed attacks, which maliciously manipulate the functional execution of Code LLMs in practice. In general, existing attacks can be divided into two main categories. (1) Backdoor attacks [46, 29, 57]. They intend to compromise the models by poisoning the training data [46, 57] or model parameters [29], such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns. They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications [29], or implementation of less secure encryption rules or protocols [46]. (2) Adversarial attacks [62, 12, 24]. They aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs.
@@
However, we reveal that existing attacks suffer from some inherent limitations, hindering their realistic threats to some extent. Specifically, backdoor attacks can achieve specific malicious objectives but require the capability of interfering the model training, e.g., modifying some training samples or even model parameters. In real-world scenarios, developers of Code LLMs or integrated applications usually obtain the models or datasets from trusted entities (e.g., renowned research institutes or companies). This mitigates the backdoor threats from the source. Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult. Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks [12, 62, 24] can hardly achieve precise malicious objectives, although they do not require intervention in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, can inevitably weaken Code LLMs’ understanding of the input. Driven by these findings, an intriguing and critical question arises: Has the application of Code LLMs already achieved a sufficient level of security in practice?
@@
Unfortunately, the answer to the above question is negative. In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger [50] and Greedy Coordinate Gradient (GCG) [63]. Please find more details in Section II.. TPIA leads to a more realistic threat by inheriting the advantageous features of both backdoor and adversarial attacks. In general, as shown in Figure 1, the TPIA attacker generates a concise non-functional perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion) , and inserts it into the victim’s code context. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations. Arguably, attackers can easily introduce particular malicious non-functional perturbations to the victim’s code context in practice. For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27]. In particular, different from supply chain attacks [41] that can be easily detected by static analysis [33] or sandboxing [9] since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
@@
Figure 1: The attack paragdim of our TPIA. A concise non-functional attacker-generated perturbation is injected into one of the project’s files. When the victim user completes the code via the Code LLM, the model will be induced by the incorporated attacker-specified code to generate an unsafe output at specified position, e.g., in edit_part() function.
@@
In general, our attack consists of two main stages. In the first stage, we construct the input and target output based on the attacker-specified objective. We represent the malicious objective as an output tuple (target position, target code), which aims to induce the Code LLM to generate the target code at the target position. We further expand the tuple into a complete task code, including a conditional code (e.g., necessary package references), context code (e.g., noises), and position code to simulate the actual attack scenario. In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective. Specifically, we obtain the base Code LLM and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code. This computation involves two important optimizations: (1) Forward reasoning enhancement: it allows the optimization direction of adversarial computation to be closer to the actual reasoning process of the Code LLM (forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. (2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the number of tokens required for a successful attack.
@@
We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages. There are a total of 31 threat cases since some of these malicious objectives apply only on C/C++. We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b, and CodeGeex2-6b. The results demonstrate that TPIA can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness. In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 535k installations in the VScode [6] extension shop) and Github Copilot [19] (developed by Github and OpenAI [4]). These applications correspond to the gray-box and black-box attack scenarios, respectively. Our TPIA achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its practical threat in the real world.
@@
    We conduct extensive experiments to evaluate TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
II Background and Related Work
II-A Code LLMs for Code Completion
@@
Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation [48, 45], and code summarization [40, 8]. The emergence of LLMs provides new alternatives to the coding tasks. By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [53], CodeLlama [44], CodeGemma [49] and CodeGeex [61]. These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast of functionally complex code [54].
@@
In this paper, we focus on code completion [34, 16, 60]. This is a paradigmatic task that involves generating new code from incomplete input using Code LLMs, which have been integrated into commercial applications [19, 2]. However, this development paradigm may encounter some potential threats. On the one hand, Code LLMs can be easily influenced by the code context, which alters the result of code generation. On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs. These threats have already occurred in the real world [7], highlighting the necessity of its security research.
II-B Backdoor and Adversarial Attacks
@@
Backdoor attacks target the training process of DNNs [20, 31, 58]. These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [26, 30, 21]. In the code domain, there are generally two strategies to achieve this, including dataset poisoning [46, 57] and model poisoning [29]. The former requires the attacker to modify the code dataset for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM for application development. However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned institutes or companies). This mitigates the backdoor threats from the source. Recently, a few works also exploited backdoor attacks for positive purposes [32, 55, 52], which are out of our scope.
@@
Adversarial attacks target the model’s inference process [11, 43, 25]. Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [59, 28, 47]. In the code domain, existing adversarial attacks [12, 62, 24] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score [42]. However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input. Accordingly, adversarial attacks are difficult to pose a serious threat to Code LLMs in practice.
II-C Prompt Injection and Indirect Prompt Injection Attacks
@@
Prompt injection (PI) attacks, are a special form of attack targeting LLMs [15, 36, 63]. A malicious LLM user aims to bypass the safety guardrails of the LLM and forces it to perform undesired responses, e.g., details of creating bomb. In the code domain, existing research [39, 37, 17] focuses on code-oriented attack objectives, such as using LLMs to create malware. In general, these attacks target the developers of models (to make their models unsafe) instead of users, which is out of the scope of this paper.
@@
Indirect prompt injection (IPI) attack attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites [22]. This attack simultaneously threatens both LLM-integrated applications (by reducing usability) and victim users (by receiving malicious responses). In the code domain, the only existing work [22] makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it does not consider stealthiness and inserts a large amount of readable comments containing the malicious intent. Arguably, the existing IPI attack does not pose a realistic threat to Code LLMs.
III Problem Statement
III-A Code Completion Task